Can you please help me with this question?I get lost.
((( 1/3(x-9)=-6 )))

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

1/3(X-9)=-6 OR X/3-9/3=-6 OR X/3-3=-6 OR X/3=-6+3 OR X/3=-3 OR X=-3*3 OR X=-9
PROOF 1/3(-9-9)=-6 OR 1/3(-18)=-6 OR -18/3=-6 OR -6=-6

Answer by stanbon(75887)

(1/3)(x-9)=-6
Use the distributive law to get:
(1/3)x-3=-6
Add 3 to both sides to get:
(1/3)x=-3
Multiply both sides by 3 to get:
x=9
Cheers,
Stan H.